The government of a developing country is implementing a new reading program in which each individual who can read spends one year teaching two others to read. At the end of the year, the original individual can stop, but each of the two people who just learned to read must now spend a year teaching two more people to read. If 1,000 people can read at the start of the program, after 10 years how many people will be able to read?
Answer by CubeyThePenguin(3113)

0 years (initial): 1000 people
1 year: 1000 + 2000
2 years: 1000 + 2000 + 4000

n years: 1000(1 + 2 + ... + 2^n) = 1000(2^n - 1)

Plug in n = 10:

1000(2^10 - 1) = 1,023,000 people